OpenStudy (raffle_snaffle):

Give a 4x4 matrix that rotates points in R^3 about the Z-axis through an angle -30*, and then translate by p=(5,-2,1)

OpenStudy (raffle_snaffle):

@Loser66

OpenStudy (loser66):

wow, you go over me. I don't know how to solve it. hehehe...

OpenStudy (raffle_snaffle):

okay thanks. Lol

OpenStudy (raffle_snaffle):

How did you obtain it?

OpenStudy (loser66):

the combination of 3x3 rotation matrix about Z-axis through an angle -30 and translation p is defined by matrices: \[\left[\begin{matrix}cos(-30)&-sin(-30)&0&5\\sin(-30)&cos(-30)&0&-2\\0&0&1&1\\0&0&0&1\end{matrix}\right]\] \[=\left[\begin{matrix}\sqrt{3}/2&1/2&0&5\\-1/2&\sqrt{3}/2&0&-2\\0&0&1&1\\0&0&0&1\end{matrix}\right]\]

OpenStudy (loser66):

hopefully it's right. :) @tkhunny

OpenStudy (loser66):

@mathmale @sourwing

OpenStudy (raffle_snaffle):

^that looks right.

OpenStudy (raffle_snaffle):

@Loser66

OpenStudy (raffle_snaffle):

the translation matrix basically shift the whole entire object in 3D. It moves it from one position to another while the rotation in the translation matrix rotates the object about the z axis.

OpenStudy (loser66):

Actually I know nothing about it. Yesterday, when you made question, I made a short research about rotation matrix and the vector translation work on it. That gave you the matrix above. Today, I forgot what I did yesterday for the part I never study before like this. hehehe... I am so sorry.!!!

OpenStudy (raffle_snaffle):

Hahaha no problem. THanks again!

OpenStudy (raffle_snaffle):

@wilks311

OpenStudy (anonymous):

You can do this, but can't do the pythag thm?! Dang dude. I'm lost lol.

OpenStudy (raffle_snaffle):

when you get this high in math you tend to forget a lot of the basics. she didn't specify which leg was 6cm lol

OpenStudy (anonymous):

Didn't matter, she specified the hypotenuse XD but man, I'm in calc. Only matrices ive done have been 2D. 3D matrix bones me lol.

OpenStudy (raffle_snaffle):

calc is easyXD

OpenStudy (raffle_snaffle):

calc 1 was the easiest of them allXD

OpenStudy (anonymous):

Yeah not saying it's hard. Just showing you that you giving me this problem is like giving a snake a pencil to use XD

OpenStudy (raffle_snaffle):

LOL